find the equation of the line determined by the given information. A. P1=(0,6);P2=(-2,3) ____________ B. P1=(4,-5);P2=(11,2) ___________ . would anyone be able to help me with these?
Answer by stanbon(75887)

find the equation of the line determined by the given information.
A. P1=(0,6);P2=(-2,3) ____________
slope = (6-3)/(0--2) = 3/2
Form: y = mx + b
You have m= 3/2 and y=6 when x = 0; solve for "b":
6 = (3/2)*0 + b
b = 6
EQUATION:
y = (3/2)x + 6
========================
B. P1=(4,-5);P2=(11,2) ___________
slope = (2--5)/(11-4) = 7/7 = 1
2 = 1*11 + b
b = -9
-----------
EQUATION:
y = x -9
=====================
Cheers,
Stan H.